OpenStudy (anonymous):

Which of the points is not a solution to 3x+2y=11

OpenStudy (anonymous):

what are the points?

OpenStudy (anonymous):

(-3,1) , (-3,10) , (1,4) (3,1)

OpenStudy (anonymous):

Just plug each of the points in and see if they equal to 11

OpenStudy (anonymous):

Can u help me FabiodaSilva

OpenStudy (anonymous):

I did not expect this to be geometry... you'r obviously looking at the formula of a circunference... I'v always been bad at geometry

OpenStudy (anonymous):

O it ok

OpenStudy (the_fizicx99):

._. this isn't geometry boi. I'll do the first one for you xD you can continue, 3x + 2y = 11 (-3,1) 3(-3) + 2(1) = 11 -9 + 2 = -7 It's not A :)

OpenStudy (anonymous):

Oh nevermind, i just realised that it's in the form of (x,y)...

OpenStudy (anonymous):

so can u help me

OpenStudy (anonymous):

Follow tHe_FiZiCx99 's example and do some work! :)

OpenStudy (anonymous):

Can anyone tell me the answer

OpenStudy (anonymous):

3x+2y=11 (-3,10) 3(-3) + 2(10) = 11 -9 + 20 = 11 20 = 11+9 20 = 20 So it's B, (-3,10)